Page MenuHomePhabricator

device asw-d-codfw.mgmt.codfw.wmnet - Inbound interface errors
Closed, ResolvedPublic

Description

Since last week Friday we are seeing some errors on xe-2/0/43 - lvs2009:enp175s0f1d1. We need to re-seat the transceiver to see if this fix the issue if not replace it.

Event Timeline

Papaul created this task.Mar 2 2020, 5:03 PM
Restricted Application added a project: Operations. · View Herald TranscriptMar 2 2020, 5:03 PM
Restricted Application added a subscriber: Aklapper. · View Herald Transcript

Mentioned in SAL (#wikimedia-operations) [2020-03-02T17:04:30Z] <vgutierrez> Stopping pybal on lvs2009 to let lvs2010 get its traffic - T246686

Papaul added a comment.Mar 2 2020, 5:47 PM

@Vgutierrez looks like re-seating the transceiver fix the problem. you can get the traffic back if we see the error again we will replace the transceiver.

Thanks.

Mentioned in SAL (#wikimedia-operations) [2020-03-02T17:50:34Z] <vgutierrez> starting pybal on lvs2009 - T246686

Papaul added a comment.Mar 2 2020, 6:04 PM

@Vgutierrez after you got the traffic back the error came back again so we will have to replace the transceiver tomorrow.

Papaul triaged this task as Medium priority.Mar 2 2020, 6:05 PM

Mentioned in SAL (#wikimedia-operations) [2020-03-03T15:45:25Z] <vgutierrez> Stopping pybal on lvs2009 to let lvs2010 get its traffic - T246686

Mentioned in SAL (#wikimedia-operations) [2020-03-03T16:05:49Z] <vgutierrez> Starting pybal on lvs2009 - T246686

Papaul closed this task as Resolved.Mar 3 2020, 4:24 PM

I think all good now on the interface . closing this task. Just replaced the transceiver on the switch side.

Laser bias current                        :  40.478 mA
    Laser output power                        :  0.6830 mW / -1.66 dBm
    Module temperature                        :  38 degrees C / 100 degrees F
    Module voltage                            :  3.2680 V
    Receiver signal average optical power     :  0.7036 mW / -1.53 dBm